OpenStudy (anonymous):

Factor the GCF first and then factor the trinomial: 2x^2 + 20x + 50 A. 2(x + 5)(x + 5) B. 2(x + 5)(x + 10) C. 2(x + 2)(x + 25) D. prime

OpenStudy (anonymous):

@AntoinetteWillman :May I help you ?:)

OpenStudy (anonymous):

Is A correct?

OpenStudy (nurali):

2x^2 + 20x + 50 2(x^2+10x+25) 2(x^2+5x+5x+25) 2(x(x+5)+5(x+5)) 2(x+5)(x+5)

OpenStudy (anonymous):

(5+x)(5+x)=25+5x+5x+x^2=10x+25+x^2 and *2 = 20x+50+2x^2 So its right !:)